# Interpreting Event-Studies from Recent Difference-in-Differences Methods

### v1

@@ 4 Practical takeaways and recommendations
The first practical takeaway from these results is that the default event-studies from CS, dCDH, and BJS should not be interpreted in the same way as traditional dynamic TWFE event-study plots. Owing to the asymmetric construction of the pre-treatment and post-treatment coefficients, a kink or jump in the plot may arise even if there is no treatment effect and parallel trends is equally violated in all periods. One should therefore not apply the typical heuristics for visual inference—e.g. looking for a discontinuity or kink at the treatment date—that may be familiar from TWFE event-studies in the non-staggered case.
@@
A natural follow-up question is to what extent the event-study plots for these alternative methods can be modified to be more comparable to conventional event-study plots. For CS/dCDH, there is a straightforward solution, which is to use “long-differences” for the pre-treatment coefficients as well as the post-treatment coefficients (i.e. always use the period before treatment as the baseline). This can be implemented, for example, in the did R and Stata packages using the options base_period = ‘‘universal’’ and long2, respectively. [^10] A blog post by Callaway (2021) provides further discussion of the differences between these two options. Similar to our ongoing example, he considers a DGP with a linear violation of parallel trends, although it only has one post-treatment period and thus does not exhibit the kink seen in our example. Nevertheless, he concludes that long-run violations of parallel trends are easier to visualize using the universal-base option, but argues that using short-differences may make it easier to detect violations of the no anticipation assumption. Using the default options in the did_multiplegt Stata package also yields comparisons using “long differences”, in contrast to the results for the R package shown above. Using these settings, the event-study estimates are numerically equivalent to the dynamic TWFE specification in the non-staggered setting considered here.
For BJS, the answers are less obvious. By design, the BJS approach uses all pre-treatment periods to compute counterfactual post-treatment outcomes—which increases efficiency under spherical errors—but this seems to lead to an inherent asymmetry between the post-treatment estimates (which assume parallel trends for all periods) and the pre-treatment estimates (which try to test this assumption). It is worth noting that the pre-treatment coefficients in the BJS event-study are still valid tests of pre-treatment parallel trends in the sense that if parallel trends holds in all pre-treatment periods, the $\beta_{r}^{BJS}$ should be zero for all $r<0$. One possible approach is then to treat the testing of pre-treatment parallel trends, and estimation of post-treatment effects assuming parallel trends in all periods as completely separate exercises. This motivates putting the BJS pre-treatment estimates on a different plot from the post-treatment estimates to avoid making misleading visual inferences. [^11] Indeed, the documentation for BJS’ Stata packages states that “A pre-trend test (for the assumptions of parallel trends and no anticipation) is a separate exercise.” Alternatively, the researcher might report an event-study for a method that treats the periods symmetrically (e.g. using the did package with the options described above) before reporting the potentially more-precise estimates from BJS that assume parallel trends in all periods.
@@
It is also worth emphasizing that the discussion here does not threaten the validity of the post-treatment event-study estimates if the appropriate parallel trends assumption holds: in this case, all of the recent methods yield interpretable treatment effect estimates even under heterogeneous treatment effects. Rather, the asymmetric construction of the pre-treatment and post-treatment event-study coefficients makes it challenging to visualize whether the estimated post-treatment coefficients could be explained by a violation of parallel trends rather than a treatment effect, at least using the conventional visual heuristics.

### v2

@@ 4 Practical takeaways and recommendations
The first practical takeaway from these results is that the default event-studies from the methods considered in Figure 1 should not be interpreted in the same way as traditional dynamic TWFE event-study plots for the purposes of evaluating parallel trends violations. Owing to the asymmetric construction of the pre-treatment and post-treatment coefficients, a kink or jump in the plot may arise even if there is no treatment effect and parallel trends is equally violated in all periods. One should therefore not apply the typical heuristics for visual inference—e.g. looking for a discontinuity or kink at the treatment date—that may be familiar from TWFE event-studies. Sensitivity analyses that compare the pre-treatment and post-treatment coefficients from these event-studies (e.g. manski_how_2017; rambachan_more_2023) may likewise yield misleading conclusions given the asymmetric construction of the pre- and post-treatment coefficients.
A natural follow-up question is: to what extent can the event-study plots for these alternative methods be modified to be more comparable to conventional event-study plots? I discuss CS and BJS in turn, followed by some discussion of additional considerations.
CS.
@@
For CS, there is a straightforward answer, which is to use “long-differences” for the pre-treatment coefficients as well as the post-treatment coefficients (i.e. always use the period before treatment as the baseline). This can be implemented, for example, in the did R and csdid Stata packages using the options base_period = ‘‘universal’’ and long2, respectively. [^9] A blog post by callaway_universal_2021 provides further discussion of the differences between these two options. Similar to our ongoing example, he considers a DGP with a linear violation of parallel trends, although it only has one post-treatment period and thus does not exhibit the kink seen in our example. Nevertheless, he concludes that long-run violations of parallel trends are easier to visualize using the universal-base option, but argues that using short-differences may make it easier to detect violations of the no anticipation assumption. Using these settings, the event-study estimates are numerically equivalent to the dynamic TWFE specification in the non-staggered setting considered here.
BJS.
For BJS, the answers are somewhat less obvious. The challenge arises from the fact that, by construction, BJS uses all of the pre-treatment data in predicting post-treatment counterfactuals for treated units (as they show, this is efficient under spherical errors). [^10] The same is true of other imputation approaches such as gardner_two-stage_2021 and liu_practical_2024. In our non-staggered setting, the post-treatment effects estimated by these approaches are equivalent to those for BJS; the methods only differ in how they create the pre-treatment coefficients. Thus, the recommendations for modifying the pre-treatment coefficients for BJS also apply to these other methods. Any analysis that does a similar imputation in the pre-treatment period will either (a) be looking at in-sample prediction fit—i.e. looking at residuals $Y_{it}-\hat{Y}_{it}$, where $Y_{it}$ was included in the training data for creating $\hat{Y}_{it}$, or (b) using predictions trained only on a sub-set of the pre-treatment data, in contrast to the predictions for post-treatment outcomes, which use all of the pre-treatment data. Nevertheless, it may be possible to construct event-study plots using imputation estimators that are more similar to the standard TWFE event-study plot than the BJS implementation. Below I discuss a few possible options, along with their pros and cons; this strikes me as an area where additional innovation in future work may also be possible.
In my ongoing non-staggered example, one natural approach to make the BJS event-study coefficients more symmetric would be to use the average of the pre-treatment periods as the reference for both the pre-treatment and post-treatment event-study coefficients. That is, we could construct
$\hat{\beta}_{r}^{BJS,new}=\left(\hat{E}[Y_{i,r+1}\mid D_{i}=1]-\hat{E}[Y_{i,r+1}\mid D_{i}=0]\right)-\left(\hat{E}[\bar{Y}_{i}^{pre}\mid D_{i}=1]-\hat{E}[\bar{Y}_{i}^{pre}\mid D_{i}=0]\right)\text{ for all $r$}.$
This would match the usual TWFE event-study up to a vertical shift of all the coefficients. That is, $\hat{\beta}^{BJS,new}_{r}=\hat{\beta}^{TWFE}_{r}+\hat{c}$ for a value $\hat{c}$ that doesn’t depend on $r$. However, it is worth noting that this construction guarantees that the modified pre-trends coefficients average to zero (i.e. $\sum_{r<0}\hat{\beta}^{BJS,new}_{r}=0$). Visual inference based on comparing the average of the event-study coefficients before and after treatment may thus be misleading, since the average in the pre-treatment period is zero by construction. Nonetheless, using this more symmetric imputation approach seems preferable to me than the default BJS approach if one wants output resembling a typical event-study, provided that the researcher keeps in mind that they should only be looking at the relative differences between the coefficients and not their averages before and after treatment. How can this approach be extended to the staggered setting? liu_cohort-anchored_2025 proposes a notion which he refers to as “block bias” which measures the pre-treatment violations of parallel trends for each adoption cohort (relative to a fixed comparison group) in a staggered setting.111111liu_cohort-anchored_2025 also discusses how these block biases can be used in conjunction with the sensitivity analysis approach of rambachan_more_2023. Applied to our non-staggered example, the pre-trends estimated by the block bias approach align with the modified estimator $\hat{\beta}_{r}^{BJS,new}$. Perhaps the most natural approach for imputation estimators in staggered settings is then to create an event-study where the pre-trends estimates are formed by taking a (weighted) average of the block biases across adoption cohorts.
A related suggestion by li_benchmarking_2025 is to adopt a “leave-one-out” approach to pre-trends estimation using imputation estimators, where for relative time $r<0$, one computes averages of the form $Y_{it}-\hat{Y}_{it}^{loo}$, where the prediction $\hat{Y}_{it}^{loo}$ does not use information from relative time $r$. This approach does not align with the standard TWFE event-study when specialized to the common timing case—the pre-trend coefficients are based on predictions using all-but-one pre-treatment period but the post-treatment coefficients are based on predictions using all the pre-treatment periods—but the differences are small when the number of pre-treatment periods is large. [^12] Another approach, which is implemented in the fect and did2s packages, reports averages of the form $Y_{it}-\hat{Y}_{it}$ for all relative time periods $r$, where $\hat{Y}_{it}$ is the same as for BJS. As discussed in more detail in Appendix Section A, this produces an event-study that shrinks the coefficients $\hat{\beta}^{BJS,new}_{r}$ towards zero in the pre-treatment period, and thus may produce event-study plots with discontinuities at the treatment date.
An alternative approach recommended by BJS is to completely separate the testing and estimation steps. [^13] The documentation for BJS’ Stata package states, “A pre-trend test (for the assumptions of parallel trends and no anticipation) is a separate exercise.” The pre-treatment coefficients in the BJS default event-study are still valid tests of pre-treatment parallel trends in the sense that if parallel trends holds in all pre-treatment periods, the $\beta_{r}^{BJS}$ should be zero for all $r<0$, and so can be used for testing the null of parallel pre-trends. [^14] See roth_pre-test_2021 and bilinski_seeking_2018 regarding pitfalls of simply looking at the significance of the pre-treatment coefficients, however. As argued in those papers, researchers should ideally also look at the magnitudes of the pre-trends estimates and their confidence intervals. However, evaluating the magnitudes of potential bias from pre-trends is challenging if the pre-treatment and post-treatment coefficients are not constructed symmetrically. However, if one adopts this approach, then I would humbly suggest putting the BJS pre-treatment estimates on a different plot from the post-treatment estimates to avoid making misleading visual inferences. If one prefers to have a typical-looking event-study plot, then a straightforward way of separating the testing and estimation steps would be for the researcher to report an event-study for a method that treats the periods completely symmetrically (e.g. using the did package with the options described above) to visually evaluate the plausibility of the parallel trends assumption, and to then subsequently report the potentially more-precise post-treatment estimates from BJS that assume parallel trends in all periods. One can also conduct sensitivity analysis for these post-treatment estimates directly using the approach in rambachan_more_2023; see liu_cohort-anchored_2025 for suggestions for implementing this approach in staggered settings in conjunction with imputation estimators like BJS.
Additional discussion.
@@
It is worth emphasizing that the discussion here does not threaten the validity of the post-treatment event-study estimates for recent methods if the appropriate parallel trends assumption holds: in this case, all of the recent methods yield interpretable treatment effect estimates even under heterogeneous treatment effects. Rather, the asymmetric construction of the pre-treatment and post-treatment event-study coefficients makes it challenging to visualize whether the estimated post-treatment coefficients could be explained by a violation of parallel trends rather than a treatment effect, at least using the conventional visual heuristics (or typical sensitivity analyses).
Finally, it is worth noting that even if a method produces an event-study plot that matches the usual TWFE event-study when specialized to the non-staggered setting, additional complications may arise when interpreting these results in staggered settings. One particularly important issue for interpretation in the staggered setting, which is discussed in detail in liu_cohort-anchored_2025, is that the composition of units that contribute to an event-study coefficient will often change across different relative time periods. Indeed, if a panel is balanced in calendar time, then with staggered treatment adoption it is necessarily imbalanced in relative time. One therefore has to be careful in interpreting aggregated event-studies since the composition of units that contribute to the coefficient for relative time $r$ will be different from that for relative time $r^{\prime}$. To assess the importance of changing composition, it will often be useful to supplement aggregated event-studies with plots of the time series of the mean outcome by adoption cohort. It may also be useful to report aggregated event-studies that enforce some balancing criteria—e.g. reporting an event-study that is restricted to units who are observed at least 3 periods before and after treatment.
Appendix A Event-study plots based on in-sample imputation
Another approach to constructing event-studies for imputation estimators, which is implemented in the fect and did2s packages, reports averages of the form $Y_{it}-\hat{Y}_{it}$ for all relative time periods $r$, where $\hat{Y}_{it}$ is the same as for BJS described above. li_benchmarking_2025 study this approach in detail, and argue that it systematically under-estimates the magnitude of the pre-trends. The issue stems from the fact that the pre-treatment outcomes for treated units are used to form $\hat{Y}_{it}$, and thus $Y_{it}-\hat{Y}_{it}$ is an in-sample prediction error for pre-treatment observations. When applied to the non-staggered setting, the results in li_benchmarking_2025 imply that this in-sample approach yields pre-treatment coefficients equal to $\frac{N_{0}}{N}\hat{\beta}^{BJS,new}_{r}$ for $r<0$, where $N_{0}$ is the number of untreated units. Thus, the pre-treatment coefficients $\hat{\beta}^{BJS,new}_{r}$ are multiplied by the factor $\frac{N_{0}}{N}<1$, which is very small when most of the units are treated. Appendix Figure 1 illustrates this attenuation by plotting the event-studies produced by fect and did2s in our simulated data: there is a clear jump in the coefficients around the treatment date.
(a) fect
(b) did2s
Appendix Figure 1: Event-studies from methods using in-sample imputation (fect and did2s)